# Towards a complete scheme of cosmological neutrino self-interactions: Collision term for a wide range of mediator masses

### v1

@@ Appendix B Collision integral reduction for $\alpha-\beta$ elastic scattering processes
$\tilde{{\cal E}}_{\ell}(z_{1})$ $=\frac{1}{q_{1}}\,\int_{\mu_{j}}^{\infty}\!dz_{2}\sqrt{z_{2}^{2}-\mu_{j}^{2}}\int_{\mu_{j}}^{z_{1}+z_{2}-\mu_{i}}\!dz_{4}\,(z_{4}^{2}-\mu_{j}^{2})\,\Psi_{S,4}\,{\cal I}_{4\ell}(z_{1},z_{2},z_{4})\,.$
@@ Appendix C On $\ell=0,1$ moments
Next, we observe that the squared invariant amplitude $|{\cal M}|^{2}$ (see App. D). is invariant under any of the following momentum exchanges
$\begin{array}[]{l l}\text{(i)}\;P_{3}\leftrightarrow P_{4}\,,&\qquad\text{(iii)}\;P_{1}\leftrightarrow P_{3}\ \text{and}\ P_{2}\leftrightarrow P_{4}\,,\\[6.0pt] \text{(ii)}\;P_{1}\leftrightarrow P_{2}\,,&\qquad\text{(iv)}\;P_{1}\leftrightarrow P_{4}\ \text{and}\ P_{2}\leftrightarrow P_{3}\,.\end{array}$ (C.3)
@@
On the other hand, any integral of the form $\int d^{3}p_{1}\,d\Omega\,A\,C^{(1)}[{\bf x},{\bf p}_{1}]$ is equal to zero if $A$ is a collisional invariant. Then
$\int d^{3}p_{1}\,d\Omega\,p_{1}\,\tilde{C}^{(1)}[{\bf k},{\bf p}_{1}]=0~.$ (C.6)
that reflects momentum conservation and also is valid for energy conservation; ${\bf p}_{i}\to E_{i}$. Similarly, by considering the (dipole) integral in Fourier space and with the use of the transformation properties under (ii) and (iii), we can write
@@
$\int d^{3}p_{1}({\bf p}_{1}\cdot\hat{\bf k})\tilde{C}^{(1)}[\mathbf{k},\mathbf{p_{1}}]=0\,$, (C.8)
@@
$\int_{0}^{\infty}dp_{1}p_{1}^{2}C_{0}=\frac{1}{4\pi}\int d^{3}p_{1}\tilde{C}^{(1)}=0\quad\text{and}\quad\int_{0}^{\infty}dp_{1}p_{1}^{3}C_{1}=\frac{1}{4\pi}\int\!d^{3}p_{1}p_{1}(\mathbf{\hat{p}_{1}}\cdot\mathbf{\hat{k}})\tilde{C}^{(1)}=0,$ (C.10)
@@ Appendix D Elastic Scattering Amplitudes for $\nu_{i}-\nu_{j}$ Via Scalar Mediator
    $\overline{\left|{\cal M}_{D}^{(ij)}\right|}^{2}=\begin{cases}|g_{ii}|^{4}\left[T_{t}^{(i)}+T_{u}^{(i)}+T_{tu}^{(i)}\right]&\text{for $m_{i}=m_{j}$}\\ |g_{ii}|^{2}|g_{jj}|^{2}T_{t}^{(ij)}&\text{for $m_{i}\neq m_{j}$}\end{cases}\,.$ (D.2)
@@
  • For Majorana-like neutrinos. We can rescale the interaction Eq. (D.1) due to Majorana fields’ properties (see Subsec. 3.1), so we can use the same couplings for our calculations, and the Lagrangian remains the form (D.1) and using the consistent Feynman rules for Majorana neutrinos (see e.g. [45]).
    $\overline{\left|{\cal M}_{M}^{(ij)}\right|}^{2}=\begin{cases}|g_{ii}|^{4}\left[T_{s}^{(i)}+T_{t}^{(i)}+T_{u}^{(i)}+T_{st}^{(i)}+T_{su}^{(i)}+T_{tu}^{(i)}\right]&\text{for $m_{i}=m_{j}$}\\ |g_{ij}|^{4}T_{s}^{(ij)}+|g_{ii}|^{2}|g_{jj}|^{2}T_{t}^{(ij)}+{\rm Re}\{g_{ij}^{2}g_{ii}^{*}g_{jj}^{*}\}T_{st}^{(ij)}&\text{for $m_{i}\neq m_{j}$}\end{cases}\,.$ (D.4)
@@ Appendix E First reduced angular integrals
$J_{1}=\int_{-1}^{1}dvH(\Delta_{v}){\cal T}_{v}^{N}\quad\text{and}\quad J_{n\ell}=\int_{-1}^{1}dvH(\Delta_{v}){\cal T}_{v}^{N}{\cal P}_{\ell}(v)\,$, (E.5)
Where $n$-labels are related directly to the labels in the momenta of the involved particles, if the final states are identical $J_{3\ell}=J_{4\ell}$.
@@
$T_{u}=\frac{(A_{2}-B_{1}^{\prime}x+B_{1}y)^{2}}{(A^{\prime}_{2}-B_{1}^{\prime}x+B_{1}y)^{2}}\,$, (E.7)
$T_{s}^{(ij)}=\begin{cases}\frac{(A_{4,ij}-B_{1}^{\prime}x)^{2}}{(A_{4,ij}^{\prime}-B_{1}^{\prime}x)^{2}}&\text{for $\mu_{i}\neq\mu_{j}$}\\ \frac{(A_{4}-B_{1}^{\prime}x)^{2}}{(A^{\prime}_{4}-B_{1}^{\prime}x)^{2}}&\text{for $\mu_{i}=\mu_{j}$}\end{cases}\,$, (E.8)
@@
$T_{tu}=\frac{A_{3}+B_{2}^{\prime}x+B_{2}y+B_{1}B_{1}^{\prime}xy-B_{1}^{2}y^{2}}{(A^{\prime}_{1}-B_{1}y)(A^{\prime}_{2}-B_{1}^{\prime}x+B_{1}y)}~,$ (E.9)
$T_{st}^{(ij)}=\begin{cases}-\frac{A_{5,ij}+B_{3,ij}^{\prime}x+B_{3,ij}y-B_{1}B_{1}^{\prime}xy}{(A_{4,ij}^{\prime}-B_{1}^{\prime}x)(A_{1,i}^{\prime}-B_{1}y)}\\ \frac{A_{8}-B_{4}^{\prime}x+B_{4}z+B_{1}^{\prime 2}x^{2}-B_{1}^{\prime}B_{1}^{\prime\prime}xz}{(A_{4,ij}^{\prime}-B_{1}^{\prime}x)(A_{7}-B_{1}^{\prime}x+B_{4}z)}\end{cases}$ (E.10)
@@
$D_{1}=\frac{a-Ab/B+A^{2}}{A+A^{\prime}}~,\quad\text{and}\quad D_{2}=\frac{a+A^{\prime}b/B+A^{\prime 2}}{A+A^{\prime}}~,$
also
where
$C_{3}=\frac{a+bA/B}{A+A^{\prime}}~,\quad\text{and}\quad C_{4}=\frac{a-bA^{\prime}/B}{A+A^{\prime}}~,$
and
$\frac{a+bw}{A-Bw}=-\frac{b}{B}+\frac{a+bA/B}{A-Bw}~.$ (E.18)
@@
The resultant Heaviside step function would restrict the limits of the next angular integrals to the region where the discriminant $\Delta_{v}$ of $S(w)$ remains positively defined. Further integrations, however, would have to proceed numerically.
@@ E.1 The $h$-functions
The previous integrals (E.19)-(E.23) fail for $w\pm q=0$, which means that there is a divergence; in our case, we can find on-shell scalar production regions or infrared divergences. In the following, we present an example of how to regularize those cases in our numerical codes.
For the channel $u$ we can write Eq. (E.7) in partial fractions as
@@
$\int_{-1}^{1}\frac{dy}{(y+q_{x})^{2}\sqrt{S(y)}}H[S(y)]$ $=-\frac{\sqrt{S(y)}}{c_{x}^{\prime\prime}(y+q_{x})}\bigg|_{\tilde{y}_{+}}^{\tilde{y}_{-}}H(\Delta_{x})-\frac{b_{x}^{\prime\prime}}{2c_{x}^{\prime\prime}}\int_{-1}^{1}\frac{dy}{(y+q_{x})\sqrt{S(y)}}H[S(y)]$ (E.28)
$=-\pi H(\Delta_{x})\frac{b_{x}^{\prime\prime}}{2c_{x}^{\prime\prime}\sqrt{-c_{x}^{\prime\prime}}}h_{1}(q_{x})\,.$
@@
We would like to remark that the contribution of the infrared divergence associated with the $u$-channel in the previous example may not need to be taken into account, as it depends on whether the relevant set of energies reaches this regime within the integration domain. Nevertheless, it is useful for our code to check whether it is necessary to include these contributions. Similarly, we have constructed the remaining functions in Eqs. (3.33), (3.34), and (3.46). Notice that this procedure is useful for excluding the resonant production and for identifying the range in which it appears.
@@ E.2 Massive mediator limit
The amplitudes for Majorana or Dirac-like neutrinos in the Heavy mediator limit $m_{\varphi}\gtrsim 10^{3}$ take the form
@@
$\begin{aligned} \Delta_{2}(y)&=q_{1}^{2}q_{2}^{2}~,\\ \Delta_{1}(y)&=q_{1}^{2}q_{2}\left[q_{3}(1-y)-2q_{2}\right]~,\\ \Delta_{0}(y)&=q_{1}^{2}\left[q_{2}^{2}-q_{2}q_{3}(1-y)+q_{3}^{2}(1-y)^{2}\right]~.\end{aligned}\quad\begin{aligned} \Delta_{2}(x)&=q_{1}^{2}q_{3}^{2}~,\\ \Delta_{1}(x)&=q_{1}^{2}q_{3}\left[q_{2}(1-x)-2q_{3}\right]~,\\ \Delta_{0}(x)&=q_{1}^{2}\left[q_{3}^{2}-q_{2}q_{3}(1-x)+q_{2}^{2}(1-x)^{2}\right]~.\end{aligned}$
@@ Appendix F Symbols
$f_{g}$ first-order, gauge-dependent contribution to the Boltzmann hierarchy equations (2.16), (2.17), (2.18) (3.35), (3.49)

### v2

@@ Appendix B Collision integral reduction for $\alpha-\beta$ elastic scattering processes
$\tilde{{\cal E}}_{\ell}(z_{1})$ $=\frac{1}{q_{1}}\,\int_{\mu_{j}}^{\infty}\!dz_{2}\sqrt{z_{2}^{2}-\mu_{j}^{2}}\int_{\mu_{j}}^{z_{1}+z_{2}-\mu_{i}}\!dz_{4}\,(z_{4}^{2}-\mu_{j}^{2})\,\Psi_{S,4}\,{\cal I}_{4\ell}(z_{1},z_{2},z_{4})\,$,
@@ Appendix C On $\ell=0,1$ moments
Next, we observe that the squared invariant amplitude $|{\cal M}|^{2}$ (see App. D) is invariant under any of the following momentum exchanges
$\begin{array}[]{l l l l}\text{(i)}\;P_{3}\leftrightarrow P_{4}\,,&\quad\text{(ii)}\;P_{1}\leftrightarrow P_{2}\,,&\text{(iii)}\;P_{1}\leftrightarrow P_{3}\ \text{and}\ P_{2}\leftrightarrow P_{4}\,,&\quad\text{(iv)}\;P_{1}\leftrightarrow P_{4}\ \text{and}\ P_{2}\leftrightarrow P_{3}\,.\end{array}$ (C.3)
@@
On the other hand, any integral of the form $\int d^{3}p_{1}\,d\Omega\,A\,C^{(1)}[{\bf x},{\bf p}_{1}]$ is equal to zero if AA is a collisional invariant. Then, the expression
$\int d^{3}p_{1}\,d\Omega\,p_{1}\,\tilde{C}^{(1)}[{\bf k},{\bf p}_{1}]=0~,$ (C.6)
reflects momentum conservation and is also valid for energy conservation; ${\bf p}_{i}\to E_{i}$. Similarly, by considering the (dipole) integral in Fourier space and with the use of the transformation properties under (ii) and (iii), we can write
@@
$\int d^{3}p_{1}({\bf p}_{1}\cdot\hat{\bf k})\tilde{C}^{(1)}[\mathbf{k},\mathbf{p_{1}}]=0\,.$ (C.8)
@@
$\int_{0}^{\infty}dp_{1}p_{1}^{2}C_{0}=\frac{1}{4\pi}\int d^{3}p_{1}\tilde{C}^{(1)}=0\quad\text{and}\quad\int_{0}^{\infty}dp_{1}p_{1}^{3}C_{1}=\frac{1}{4\pi}\int\!d^{3}p_{1}p_{1}(\mathbf{\hat{p}_{1}}\cdot\mathbf{\hat{k}})\tilde{C}^{(1)}=0.$ (C.10)
@@ Appendix D Elastic Scattering Amplitudes for $\nu_{i}-\nu_{j}$ Via Scalar Mediator
    $\overline{\left|{\cal M}_{D}^{(ij)}\right|}^{2}=\begin{cases}|g_{ii}|^{4}\left[T_{t}^{(i)}+T_{u}^{(i)}+T_{tu}^{(i)}\right]&\text{for $m_{i}=m_{j}$}\\ |g_{ii}|^{2}|g_{jj}|^{2}T_{t}^{(ij)}&\text{for $m_{i}\neq m_{j}$}\end{cases}\,$, (D.2)
@@
  • For Majorana-like neutrinos. We can rescale the interaction Eq. (D.1) due to Majorana fields’ properties (see Subsec. 3.1), so we can use the same couplings for our calculations, the Lagrangian remains the form (D.1) and using the consistent Feynman rules for Majorana neutrinos (see, e.g., Ref. [46]).
    $\overline{\left|{\cal M}_{M}^{(ij)}\right|}^{2}=\begin{cases}|g_{ii}|^{4}\left[T_{s}^{(i)}+T_{t}^{(i)}+T_{u}^{(i)}+T_{st}^{(i)}+T_{su}^{(i)}+T_{tu}^{(i)}\right]&\text{for $m_{i}=m_{j}$}\\ |g_{ij}|^{4}T_{s}^{(ij)}+|g_{ii}|^{2}|g_{jj}|^{2}T_{t}^{(ij)}+{\rm Re}\{g_{ij}^{2}g_{ii}^{*}g_{jj}^{*}\}T_{st}^{(ij)}&\text{for $m_{i}\neq m_{j}$}\end{cases}\,$, (D.4)
@@ Appendix E First reduced angular integrals
$J_{1}=\int_{-1}^{1}dvH(\Delta_{v}){\cal T}_{v}^{N}\quad\text{and}\quad J_{n\ell}=\int_{-1}^{1}dvH(\Delta_{v}){\cal T}_{v}^{N}{\cal P}_{\ell}(v)\,.$ (E.5)
Where $n$-labels are related directly to the labels in the momenta of the involved particles, if the final states are identical, $J_{3\ell}=J_{4\ell}$.
@@
$T_{u}=\frac{(A_{2}-B_{1}^{\prime}x+B_{1}y)^{2}}{(A^{\prime}_{2}-B_{1}^{\prime}x+B_{1}y)^{2}}\,,\,T_{s}^{(ij)}=\begin{cases}\frac{(A_{4,ij}-B_{1}^{\prime}x)^{2}}{(A_{4,ij}^{\prime}-B_{1}^{\prime}x)^{2}}&\text{for $\mu_{i}\neq\mu_{j}$}\\ \frac{(A_{4}-B_{1}^{\prime}x)^{2}}{(A^{\prime}_{4}-B_{1}^{\prime}x)^{2}}&\text{for $\mu_{i}=\mu_{j}$}\end{cases},$ (E.7)
@@
$T_{tu}=\frac{A_{3}+B_{2}^{\prime}x+B_{2}y+B_{1}B_{1}^{\prime}xy-B_{1}^{2}y^{2}}{(A^{\prime}_{1}-B_{1}y)(A^{\prime}_{2}-B_{1}^{\prime}x+B_{1}y)}\,,\,T_{st}^{(ij)}=\begin{cases}-\frac{A_{5,ij}+B_{3,ij}^{\prime}x+B_{3,ij}y-B_{1}B_{1}^{\prime}xy}{(A_{4,ij}^{\prime}-B_{1}^{\prime}x)(A_{1,i}^{\prime}-B_{1}y)}\\ \frac{A_{8}-B_{4}^{\prime}x+B_{4}z+B_{1}^{\prime 2}x^{2}-B_{1}^{\prime}B_{1}^{\prime\prime}xz}{(A_{4,ij}^{\prime}-B_{1}^{\prime}x)(A_{7}-B_{1}^{\prime}x+B_{4}z)}\end{cases},$ (E.8)
@@
$D_{1}=\frac{a-Ab/B+A^{2}}{A+A^{\prime}}\,,\,D_{2}=\frac{a+A^{\prime}b/B+A^{\prime 2}}{A+A^{\prime}}\,,\,C_{3}=\frac{a+bA/B}{A+A^{\prime}}\quad\text{and}\quad C_{4}=\frac{a-bA^{\prime}/B}{A+A^{\prime}}\,.$
@@
The resulting Heaviside function restricts the angular integrals to regions where the discriminant $\Delta_{v}$ of $S(w)$ is positive; further integrations must be performed numerically.
@@ E.1 The $h$-functions
The integrals (E.17)-(E.21) fail for $w\pm q=0$, which corresponds to either on-shell scalar production or an infrared divergence. In the following, we illustrate how to regularize these cases in our numerical implementation.
For the $u$-channel we can write Eq. (E.7) in partial fractions as
@@
$\int_{-1}^{1}\frac{dy}{(y+q_{x})^{2}\sqrt{S(y)}}H[S(y)]$ $=-\frac{\sqrt{S(y)}}{c_{x}^{\prime\prime}(y+q_{x})}\bigg|_{\tilde{y}_{+}}^{\tilde{y}_{-}}H(\Delta_{x})-\frac{b_{x}^{\prime\prime}}{2c_{x}^{\prime\prime}}\int_{-1}^{1}\frac{dy}{(y+q_{x})\sqrt{S(y)}}H[S(y)]$
$=-\pi H(\Delta_{x})\frac{b_{x}^{\prime\prime}}{2c_{x}^{\prime\prime}\sqrt{-c_{x}^{\prime\prime}}}h_{1}(q_{x})\,$, (E.26)
@@
We note that the infrared divergence in the $u$-channel of the previous example may be neglected, as it depends on whether the relevant energies lie within the integration domain. Nevertheless, it is useful for our code to check whether it is necessary to include these contributions. Similarly, we have constructed the remaining functions in Eqs. (3.33), (3.34), and (3.46). Notice that this procedure is useful for excluding the resonant production and for identifying the range in which it appears.
@@ E.2 Massive mediator limit
The amplitudes for Majorana or Dirac-like neutrinos in the HML take the form
@@
$\begin{aligned} \Delta_{2}(y)&=q_{1}^{2}q_{2}^{2}~,\\ \Delta_{1}(y)&=q_{1}^{2}q_{2}\left[q_{3}(1-y)-2q_{2}\right]~,\\ \Delta_{0}(y)&=q_{1}^{2}\left[q_{2}^{2}-q_{2}q_{3}(1-y)+q_{3}^{2}(1-y)^{2}\right]~.\end{aligned}\quad\begin{aligned} \Delta_{2}(x)&=q_{1}^{2}q_{3}^{2}~,\\ \Delta_{1}(x)&=q_{1}^{2}q_{3}\left[q_{2}(1-x)-2q_{3}\right]~,\\ \Delta_{0}(x)&=q_{1}^{2}\left[q_{3}^{2}-q_{2}q_{3}(1-x)+q_{2}^{2}(1-x)^{2}\right]~,\end{aligned}$
@@ Appendix F Symbols
$f_{g,\ell}$ first-order, gauge-dependent contribution to the Boltzmann hierarchy equations (2.16), (2.17), (2.18) (3.35), (3.49)